OpenStudy (anonymous):

can somebdoy check if i got some questions right please 1. √48 i got 4√3 2. √3/25 i got (31/2)/5 3. √25/16 i got 5/2 4.√ 2^2/b^5 i got (a/b^2(b^1/2) 5. √a/b√a^2/b^2 i got (a/b)(a/b)^1/2

OpenStudy (anonymous):

Will you explain what you get in Second one...???

OpenStudy (anonymous):

You have first one correctly.. In third one, what is the square root of 16???

OpenStudy (anonymous):

second(sqrt(3)/5)

OpenStudy (anonymous):

third 5/4

OpenStudy (anonymous):

fourth 2.sqrt(b)/b³

OpenStudy (anonymous):

can u guys explain please

OpenStudy (anonymous):

second one i got 31/2/5

OpenStudy (anonymous):

why 31?

OpenStudy (anonymous):

sq of 16 is 4

OpenStudy (anonymous):

i dont know im very confused :(

OpenStudy (anonymous):

yes ,but 31?

OpenStudy (anonymous):

how do you get 31?

OpenStudy (anonymous):

why not 17 lol

OpenStudy (anonymous):

i dont know how to do these i dont know how i got them

OpenStudy (anonymous):

May be the second one you got will be: \[(3)^{1/2}/5\] Is it??? If it is, then it is correct...

OpenStudy (anonymous):

yesssss i did get that

OpenStudy (anonymous):

Then you are right...

OpenStudy (anonymous):

what about the otherss

OpenStudy (anonymous):

Please explain how you did 4th one...

OpenStudy (anonymous):

its not 2^2 its a^2

OpenStudy (anonymous):

the 4th one im not really sure

OpenStudy (anonymous):

\[\sqrt{a^2/b^5} = a/(b^2.b^{1/2})\] Got it or not???

OpenStudy (anonymous):

yes got it

OpenStudy (anonymous):

5th One: \[\sqrt{a/b}.\sqrt{a^2/b^2} = (a/b).(a/b)^{1/2}\] Got or not??

OpenStudy (anonymous):

waittt the one you did before this do you just leave the answer like that or simplify ahead?

OpenStudy (anonymous):

and this one too just leave it like that

OpenStudy (anonymous):

You can leave it in the form as I have written also...

OpenStudy (anonymous):

Or you may write it as: \[(a/b^2).(1/b^{1/2})\]

OpenStudy (anonymous):

for the first one or the second one? it says simplify so i guess the most simplest form sorry one more question where are you getting the 1/2 from

OpenStudy (anonymous):

See, \[\sqrt{}\] can also be written as : 1/2 For example: \[\sqrt{x}\] can be written as: \[x^{1/2}\] Getting nisa or not??

OpenStudy (anonymous):

ohh so thats why i see that so much

OpenStudy (anonymous):

You have done first one very correctly...

OpenStudy (anonymous):

Second One: \[\sqrt{3/25} = \sqrt{3}/\sqrt{25} = \sqrt{3}/5\]

OpenStudy (anonymous):

my cousin helped me with that one i didnt understand it

OpenStudy (anonymous):

The Second Answer can also be written as: \[3^{1/2}/5\]

OpenStudy (anonymous):

First one I draw for you: